# PSDiffusion: Harmonized Multi-Layer Image Generation via Layout and Appearance Alignment

### v1

@@
Abstract.
Diffusion models have made remarkable advancements in generating high-quality images from textual descriptions. Recent works like LayerDiffuse have extended the previous single-layer, unified image generation paradigm to transparent image layer generation. However, existing multi-layer generation methods fail to handle the interactions among multiple layers such as rational global layout, physics-plausible contact and visual effects like shadows and reflections while maintaining high alpha quality. To solve this problem, we propose PSDiffusion, a unified diffusion framework for simultaneous multi-layer text-to-image generation. Our model can automatically generate multi-layer images with one RGB background and multiple RGBA foregrounds through a single feed-forward process. Unlike existing methods that combine multiple tools for post-decomposition or generate layers sequentially and separately, our method introduces a global-layer interactive mechanism that generates layered-images concurrently and collaboratively, ensuring not only high quality and completeness for each layer, but also spatial and visual interactions among layers for global coherence.
Multi-Modal Generation, Diffusion Model, Layered Images
copyright: noneccs: Computing methodologies Image and video acquisition
@@ Abstract.
Figure 1. Given a complex prompt, our model, PSDiffusion, produces plausible per-layer layouts for each instance and achieves harmonious inter-layer interactions. Compared to the state-of-the-art (Zhang and Agrawala, 2024), PSDiffusion generates more consistent appearances with realistic visual effects (e.g., shadows for the boat and swan), without manual refinement. [Best viewed with zoom-in]
1. Introduction
@@
Diffusion-based generative models such as Stable Diffusion (Rombach et al., 2022) have revolutionized image synthesis by generating high-fidelity, diverse visuals from textual prompts, but they are limited to producing images as unified, single-layer outputs which are not easy to edit. In contrast, images with layered representations provide human designers with great advantages in image manipulations, such as precise editing through element isolation, compositional flexibility via layer recombination, and collaborative iteration through asset sharing.
The evolution of transparent image layer generation began with post-processing solutions applied to standard RGB outputs. Early works like PP-Matting (Chen et al., 2022) and Matting-Anything (Li et al., 2024) pioneered trimap-free alpha estimation for single transparent layer extraction. Subsequent pipelines such as MuLAn  (Tudosiu et al., 2024), PACO (Liu et al., 2024a) and LayeringDiff (Kang et al., 2025) established systematic frameworks for decomposing images into editable layers through detection, segmentation and completion. While operationally sound, these multi-stage methods tend to accumulate errors due to their sequential processing pipelines, primarily manifested in two aspects: challenges in generating high-quality alpha channels for foreground layers and inconsistencies arising from amodal completion operations.
@@
This limitation catalyzed the emergence of direct RGBA synthesis paradigms. Text2Layer (Zhang et al., 2023b) marked a pivotal shift by jointly generating foreground/background layers through coordinated latent space optimization in Stable Diffusion. LayerDiffuse’s (Zhang and Agrawala, 2024) ”latent transparency” mechanism accelerated the field, which enabled pretrained LDMs to produce high-fidelity single transparent layers, and Alfie’s (Quattrini et al., [n. d.]) diffusion transformer architecture that dynamically derived alpha channels from attention maps. Building on transparent single-layer image generation, recent works have made strides toward multi-layer synthesis. LayerDiffuse (Zhang and Agrawala, 2024) itself extended its architecture for sequential background-to-foreground generation . Fontanella et al. (Fontanella et al., [n. d.]) explored independent multiple foreground layer generation and post-composition. LayerDiff (Huang et al., 2024) attempted to generate multi-layer images simultaneously, but it can only generate isolated, non-overlapping layers with binary masks. These paradigms mainly depend on straightforward layer stacking instead of considering multiple layers in a holistic manner. Therefore, they tend to neglect the interactions among layers, including comprehensive layout, physics-plausible contacts and visual effects like shadows and reflections, while these are essential to cohesive multi-layer generation and editing.
@@
Another critical challenge underlying this mult-layer image generation stems from the scarcity of high-quality multi-layer RGBA image datasets. While text-image pairs abound in datasets like LAION (Schuhmann et al., 2021), commercially restricted access to premium transparent assets severely limits usable multi-layer RGBA data. Existing open-source multi-layer datasets exhibit complementary weaknesses. MAGICK (Burgert et al., 2024) and Multi-layer Dataset (Zhang and Agrawala, 2024) provide visually coherent layers but are constrained to 1-2 layers per composition, whereas MuLAn (Tudosiu et al., 2024) supports 2-6 layers per image but suffers from poor alpha-matte quality because it relies on automated post-processing pipelines.
@@
To overcome these dataset limitations, we propose the Inter-Layer Dataset, which consists of 30K meticulously curated multi-layer images, each containing 3-6 layers with artist-grade alpha mattes and rich layer interactions. Recognizing that automated synthesis and post-editing inherently fail to guarantee precise alpha channels and contextually coherent inpainting, we engaged professional designers to construct this dataset through a human-in-the-loop workflow. Designers curated layers from diverse sources including web repositories, BG20K (Li et al., 2022), MAGICK (Burgert et al., 2024), and LayerDiffuse-generated transparent assets (Zhang and Agrawala, 2024), and then applied meticulous edits to ensure natural layer interactions. Each composition integrates spatial layout optimization, physics-aware interaction and global stylistic harmony.
@@
Although existing pretrained text-to-image generation models can only generate unified single layer images, given a global text prompt describing the background, multiple foregrounds and their interactions, they did demonstrate prominent capability to spontaneously arrange the layout of multiple elements, ensure natural interactions between constituent subjects, and maintain global harmony and consistency in generated images. We believe that these pretrained text-to-image generation can provide interaction priors for multi-layer generation. Building on this fact, we propose PSDiffusion, a unified diffusion framework that utilizes the interaction priors from pretrained diffusion models, to simultaneously generate multi-layer images with coherent global structure and realistic layer interactions. We design an attention-based global-layer interactive mechanism to model the layer interactions paradigm. For one thing, to achieve a harmonious global layout for foreground layers, we develop a layer cross-attention reweighting module, where we extract the cross-attention map from global layer to guide the position of foreground layers. For another, to facilitate content coherence and interaction among different layers, and avoid layer entanglement, we develop a partial joint self-attention module, where we encourage content sharing from global layer across the foreground and background layers.
@@
2. Related Works
2.1. Text-to-Image Generation
@@
Diffusion models have made significant advancements in image generation, particularly in text-guided image synthesis, leading to remarkable improvements in both the quality of generated images and the ability to follow complex prompts. The development of large-scale diffusion models trained on extensive datasets of paired text and images has established new benchmarks in this domain, with notable examples including Stable Diffusion (Rombach et al., 2022), SDXL (Podell et al., 2024), Imagen (Saharia et al., 2022), and DALL-E 3 (Betker et al., 2023). Furthermore, recent advancements in network architecture, especially with Diffusion Transformer networks (DiT) like Stable Diffusion 3 (Esser et al., 2024) and Flux (Labs, 2024) , have enhanced image fidelity and diversity. Moreover, other research efforts focus on integrating various control conditions into pre-trained image foundation models to enhance their controllability. For instance, Gligen (Li et al., 2023), ControlNet (Zhang et al., 2023a) and IP-Adapter (Ye et al., 2023) offer greater flexibility in controllable image generation, allowing users to precisely manipulate aspects such as layout, depth, human pose, and style. However, these models can only generate a single RGB image during a single forward pass, rather than producing layered images, which are widely used formats that enable users to edit specific parts of an image without affecting other areas.
2.2. Layered Image Generation
@@
Image Layer Extraction with Post-RGB Process. Multi-layer image generation can be achieved primarily through two different paths. The first path, termed Post-RGB Process, operates through a sequential pipeline: (1) initial RGB image synthesis via diffusion models (Rombach et al., 2022); (2) foreground extraction using open-vocabulary object detection and segmentation tools (e.g. GroundingDINO (Liu et al., 2024b), SAM (Kirillov et al., 2023), Matting-Anything (Li et al., 2024)); (3) background inpainting and foreground completion for occluded areas to derive multiple RGBA layers (Rombach et al., 2022; Podell et al., 2024; Yu et al., 2023; Ao et al., 2024). Along this way, Tudosiuet et al. (Tudosiu et al., 2024) developed a training-free pipeline, which decomposes a monocular RGB image into a stack of RGBA layers comprising background and isolated instances, and released the MuLAn dataset. LayeringDiff  (Kang et al., 2025) begins by generating using an off-the-shelf image generative model, followed by disassembling the image into its constituent foreground and background layers. Yang et al. proposed LayerDecomp (Yang et al., 2024), a generative training framework that leverages both simulated and real-world data to enable robust layer-wise decomposition with accurate visual effect representation through a consistency loss, facilitating high-quality, training-free downstream editing. Generative Omnimatte (Lee et al., 2024) extends this idea to decompose video into layers and demonstrate high-quality decompositions and editing results. Despite the operational validity of these methods, the multi-stage post-processing approach imposes a nontrivial computational redundancy and risks compounding errors through iterative operations—a limitation that frequently manifests itself as incoherent visual semantics or fragmented stylistic attributes.
@@
Direct Transparent Image Layer Generation. RGBA direct generation focuses on synthesizing image layers directly, bypassing the traditional pipeline of generating a full image first and then extracting transparent layers. The earliest work, Text2Layer (Zhang et al., 2023b), extends the Stable Diffusion architecture into a dual-layer generation framework. It jointly synthesizes a background scene and semantically aligned foreground elements through coordinated latent space optimization. Subsequently, LayerDiffuse (Zhang and Agrawala, 2024) introduces ”latent transparency”, enabling large-scale pretrained latent diffusion models to generate either single transparent images. This approach has gained a reputation for its high-quality single-layer transparency generation. In parallel, Alfie (Quattrini et al., [n. d.]) and Diffumatting (Hu et al., 2024a) also develop different pipelines to generate single-layer RGBA images.
@@
Building on single-layer generation capabilities, multi-layer synthesis is achieved through either sequential or separate generation paradigms. Specifically, LayerDiffuse (Zhang and Agrawala, 2024) adopts a sequential approach, proposing a background-conditioned transparent layer generation that iteratively synthesizes layers from background to foreground. Fontanella et al. (Fontanella et al., [n. d.]) and LayerFusion (Dalva et al., 2024) employ a separate paradigm, where individual foreground layers are generated independently and then composited.
While methods of multi-layer generation based on single-layer generation have advanced RGBA synthesis, their extension suffers from insufficient layer interaction, leading to critical practical limitations. Common inter-layer contamination issues include rational global layout, physics-plausible contacts and visual effects like shadows and reflections. These problems are not easily resolvable through additional post-editing, particularly for the latter two issues.
@@
Recent unified text-guided multi-layer frameworks have emerged and shown promising results. Specifically, LayerDiff (Huang et al., 2024) proposes a simultaneous multi-layer synthesis framework using a unified diffusion model. However, it produces low-quality outputs limited to binary masks instead of full alpha channels, failing to handle overlapping foregrounds and visual effects which are essential for layer editing. ART (Pu et al., 2025) introduces an Anonymous Region Layout Planner, aligning visual tokens with text tokens to directly synthesize spatially coherent multi-layer images. While innovative, it relies on precise bounding boxes provided by users or generated by LLMs. Additionally, its anonymous layout mechanism results in layer entanglement issue, where semantically distinct elements (e.g., foreground characters and background props) are incorrectly merged into one layer, adding unnecessary complexity.
In contrast, our proposed PSDiffusion is a unified diffusion framework for simultaneous text to multi-layer image generation, ensuring not only high quality and completeness for each layer, but also spatial and semantic interaction among layers for global coherence. We adopt global prompt and layer-specific prompts as input to facilitate customized element separation and avoid layer entanglement.
@@
Figure 2. Architecture of PSDiffusion. Our framework processes multi-layer compositions through three key components: (1) A transparent VAE encoder preserving alpha channels; (2) Layer cross-attention reweighting for layout plausibility; (3) Partial joint self-attention for inter-layer context modeling.
3. Method
3.1. Problem Formulation
@@
A multi-layer composable image is defined as a structure containing a background layer $B=I_{c}^{B}\in\mathbb{R}^{H\times W\times 3}$ (RGB format) and $K$ foreground layers $\{F_{k}\}_{k=1}^{K}$ in RGBA format, where each $F_{k}=\left(I_{c}^{F_{k}},I_{\alpha}^{F_{k}}\right)$ consists of color channels $I_{c}^{F_{k}}\in\mathbb{R}^{H\times W\times 3}$ and an alpha channel $I_{\alpha}^{F_{k}}\in\mathbb{R}^{H\times W\times 1}$. Following the premultiplied alpha convention, each foreground is preprocessed as $F^{\prime}_{k}=I_{c}^{F_{k}}\ast I_{\alpha}^{F_{k}}$ through element-wise multiplication.
@@
(1) $I_{k}=\left(1-I_{\alpha}^{F_{k}}\right)\cdot I_{k-1}+I_{\alpha}^{F_{k}}\ast I_{c}^{F_{k}},\ k\in[K]$
where $F_{k}$ denotes the $k$-th foreground. After $K$ iterations, the final image $I_{K}$ composes into
(2) $I_{K}=\underbrace{\left(\prod_{k=1}^{K}\left(1-I_{\alpha}^{F_{k}}\right)\right)\ast I_{c}^{B}}_{\text{\scriptsize Background visibility attenuated}}+\underbrace{\sum_{k=1}^{K}\left(\prod_{j=k+1}^{K}\left(1-I_{\alpha}^{F_{j}}\right)\ast I_{\alpha}^{F_{k}}\ast I_{c}^{F_{k}}\right)}_{\text{\scriptsize Foreground contributions with occlusion}}$
@@
This mathematical formulation explicitly captures two fundamental mechanisms: (1) The multiplicative attenuation of background visibility through successive transparency operations; (2) The nested modulation of foreground layer contributions based on depth-ordered occlusions.
3.2. PSDiffusion

### v2

@@
Abstract
Transparent image layer generation plays a significant role in digital art and design workflows. Existing methods typically decompose transparent layers from a single RGB image using a set of tools or generate multiple transparent layers sequentially. Despite some promising results, these methods often limit their ability to model global layout, physically plausible interactions, and visual effects such as shadows and reflections with high alpha quality due to limited shared global context among layers. To address this issue, we propose PSDiffusion, a unified diffusion framework that leverages image composition priors from pre-trained image diffusion model for simultaneous multi-layer text-to-image generation. Specifically, our method introduces a global layer interaction mechanism to generate layered images collaboratively, ensuring both individual layer quality and coherent spatial and visual relationships across layers. We include extensive experiments on benchmark datasets to demonstrate that PSDiffusion is able to outperform existing methods in generating multi-layer images with plausible structure and enhanced visual fidelity.
@@ Abstract
Figure 1: Given a complex prompt, our model, PSDiffusion, produces plausible per-layer layouts for each instance and achieves harmonious inter-layer interactions. Compared to the state-of-the-art [zhang2024transparent], PSDiffusion generates layered images with more plausible spatial arrangement and more consistent appearances with realistic visual effects (e.g., reflections for the boat and swan), without manual spatial refinement. [Best viewed with zoom-in]
1 Introduction
@@
Diffusion models [rombach2022high, dalle3, imagen] have revolutionized image synthesis by generating high-fidelity, diverse visuals from textual prompts, but they are limited to producing images as unified, single-layer outputs which are not easy to edit. In contrast, images with layered representations provide human designers with great advantages in image manipulations, such as precise editing through element isolation, compositional flexibility via layer recombination, and collaborative iteration through asset sharing.
@@
This limitation of pretrained generative models catalyzed the emergence of RGBA image layer synthesis. Text2Layer [Text2Layer] explored generating image layers through segmentation guidance. LayerDiffuse’s [zhang2024transparent] “latent transparency” mechanism accelerated the field, which enabled pretrained LDMs to produce high-fidelity single transparent layers. Building on transparent single-layer image generation, recent works have made strides toward multi-layer synthesis. LayerDiffuse [zhang2024transparent] itself extended its architecture for sequential background-to-foreground generation . Fontanella et al[fontanellagenerating] explored separate multiple foreground layers generation and post-composition. LayerDiff [huang2024layerdiff] attempted to generate multi-layer images simultaneously, but it can only generate isolated, non-overlapping layers with binary masks. These paradigms mainly depend on straightforward layer stacking instead of considering multiple layers in a holistic manner. Therefore, they tend to neglect the interactions among layers, including comprehensive layout, physics-plausible contacts and visual effects like shadows and reflections, while these are essential to cohesive multi-layer generation and editing.
@@
Another critical challenge underlying this mult-layer image generation stems from the scarcity of high-quality multi-layer RGBA image datasets. While text-image pairs abound in datasets like LAION [schuhmann2021laion400mopendatasetclipfiltered], commercially restricted access to premium transparent assets severely limits usable multi-layer RGBA data. Existing open-source multi-layer datasets exhibit complementary weaknesses. MAGICK [Burgert_2024_CVPR] and Multi-layer Dataset [zhang2024transparent] provide high quality layers but are constrained to 1-2 layers per composition, whereas MuLAn [tudosiu2024mulan] supports 2-6 layers per image but suffers from poor alpha-matte quality because it relies on automated post-processing pipelines.
@@
To overcome these dataset limitations, we propose the Inter-Layer Dataset, which consists of 30K meticulously curated multi-layer images, each containing 3-6 layers with artist-grade alpha mattes and rich layer interactions. Recognizing that automated synthesis and post-editing inherently fail to guarantee precise alpha channels and contextually coherent inpainting, we engaged about 35 professional designers to construct this dataset through a human-in-the-loop image composition workflow. Each composition operation involves spatial layout optimization, physics-aware layer interactions, and global stylistic consistency.
@@
Although existing pretrained text-to-image generation models can only generate unified single layer images, given a global text prompt describing the background, multiple foregrounds and their interactions, they did demonstrate prominent capability to spontaneously arrange the layout of multiple elements, ensure natural interactions between constituent subjects, and maintain global harmony and consistency in generated images. We believe that these pretrained text-to-image generation models can provide interaction priors for multi-layer generation. Building on this fact, we propose PSDiffusion, a unified diffusion framework that utilizes the interaction priors from pretrained diffusion models, to simultaneously generate multi-layer images with coherent global structure and realistic layer interactions. We design an attention-based global-layer interactive mechanism to model the layer interactions paradigm. For one thing, to achieve a harmonious global layout for foreground layers, we develop a layer cross-attention reweighting module, where we extract the cross-attention map from the global layer to guide the position of foreground layers. For another, to facilitate content coherence and interaction among different layers, and avoid layer entanglement, we develop a partial joint self-attention module, where we encourage content sharing from global layer across the foreground and background layers.
@@
Figure 2: Visual results of PSDiffusion. Through the design of Layer Cross-Attention Reweighting Module and Partial Joint Self-Attention Mechanism, PSDiffusion synthesize multi-layer transparent images with plausible layout arrangements and harmonious appearance.
2 Related Work
2.1 Text-to-Image Generation
@@
Diffusion models have made significant advancements in image generation, particularly in text-guided image synthesis, leading to remarkable improvements in both the quality of generated images and the ability to follow complex prompts. The development of large-scale diffusion models trained on extensive datasets of paired text and images has established new benchmarks in this domain, with notable examples including Stable Diffusion [rombach2022high], SDXL [podell2023sdxl], Imagen [imagen], and DALL-E 3 [dalle3]. Furthermore, recent advancements in network architecture, especially with Diffusion Transformer networks (DiT) like Stable Diffusion 3 [esser2024scaling] and Flux [flux2024] , have enhanced image fidelity and diversity. Moreover, other research efforts focus on integrating various control conditions into pre-trained image foundation models to enhance their controllability. For instance, Gligen [li2023gligen], ControlNet [zhang2023control] and IP-Adapter [ye2023ip] offer greater flexibility in controllable image generation, allowing users to precisely manipulate aspects such as layout, depth, human pose, and style. However, these models can only generate a single RGB image during a single forward pass, rather than producing layered images, which are widely used formats that enable users to edit specific parts of an image without affecting other areas.
2.2 Layered Image Generation
@@
Image Layer Extraction with Post-RGB Process. Multi-layer image generation can be achieved primarily through two different paths. The first path, termed Post-RGB Process, operates through a sequential pipeline: (1) initial RGB image synthesis via diffusion models [rombach2022high]; (2) foreground extraction using open-vocabulary object detection and segmentation tools (e.g. GroundingDINO [liu2024grounding], SAM [Kirillov2023Segment], Matting-Anything [li2023matting]); (3) background inpainting and foreground completion for occluded areas to derive multiple RGBA layers [rombach2022high, podell2023sdxl, yu2023inpaint, ao2024open]. Along this way, Tudosiuet et al[tudosiu2024mulan] developed a training-free pipeline, which decomposes a monocular RGB image into a stack of RGBA layers comprising background and isolated instances, and released the MuLAn dataset. LayeringDiff [kang2025layeringdiff] begins by generating using an off-the-shelf image generative model, followed by disassembling the image into its constituent foreground and background layers. Yang et al. proposed LayerDecomp [yang_Layer_Decomposition], a generative training framework that leverages both simulated and real-world data to enable layer-wise post-decomposition with visual effect representation. Despite the operational validity of these methods, the multi-stage post-processing approach imposes a nontrivial computational redundancy and risks compounding errors through iterative operations—a limitation that frequently manifests itself as incoherent visual semantics or poor alpha quality.
@@
Direct Transparent Image Layer Generation. RGBA direct generation focuses on synthesizing image layers directly, bypassing the traditional pipeline of generating a full image first and then extracting transparent layers. The earliest work, Text2Layer [Text2Layer], extends the Stable Diffusion architecture into a dual-layer generation framework. It jointly synthesizes a background scene and semantically aligned foreground elements through coordinated latent space optimization. Subsequently, LayerDiffuse [zhang2024transparent] introduces ”latent transparency”, enabling large-scale pretrained latent diffusion models to generate either single transparent images. This approach has gained a reputation for its high-quality single-layer transparency generation. In parallel, Alfie [quattrini2024alfie] and Diffumatting [hu2024diffumatting] also develop different pipelines to generate single-layer RGBA images.
@@
Building on single-layer generation capabilities, multi-layer synthesis is achieved through either sequential or separate generation paradigms. Specifically, LayerDiffuse [zhang2024transparent] adopts a sequential approach, proposing a background-conditioned transparent layer generation that iteratively synthesizes layers from background to foreground. Fontanella et al[fontanellagenerating] and LayerFusion [dalva2024layerfusion] employ a separate paradigm, where individual foreground layers are generated independently and then composited.
@@
Recently unified text-guided multi-layer generation has attracted increasing attention. Specifically, LayerDiff [huang2024layerdiff] proposes a simultaneous multi-layer synthesis framework using a unified diffusion model with 3D convolutions. However, it produces low-quality outputs limited to binary masks instead of full alpha channels, failing to handle overlapping foregrounds and visual effects which are essential for layer editing. ART [pu2025art] introduces an Anonymous Region Layout Planner, aligning visual tokens with text tokens to directly synthesize spatially coherent multi-layer images. However, it relies on precise bounding boxes provided by users or generated by LLMs and its anonymous layout mechanism results in layer entanglement issue, where semantically distinct elements (e.g., foreground characters and background props) are incorrectly merged into one layer, and cause issues on alpha quality.
@@
Figure 3: Architecture of PSDiffusion. Our framework processes multi-layer compositions through three key components: (1) A transparent VAE encoder preserving alpha channels; (2) Layer cross-attention reweighting for layout plausibility; (3) Partial joint self-attention for inter-layer context modeling.
3 Methodology
3.1 Problem Formulation
@@
A multi-layer composable image is defined as a structure containing a background layer $B=I_{c}^{B}\in\mathbb{R}^{H\times W\times 3}$ in RGB format and $K$ foreground layers $\{F_{k}\}_{k=1}^{K}$ in RGBA format, where each $F_{k}=\left(I_{c}^{F_{k}},I_{\alpha}^{F_{k}}\right)$ consists of color channels $I_{c}^{F_{k}}\in\mathbb{R}^{H\times W\times 3}$ and an alpha channel $I_{\alpha}^{F_{k}}\in\mathbb{R}^{H\times W\times 1}$. Following the premultiplied alpha convention, each foreground is preprocessed as $F^{\prime}_{k}=I_{c}^{F_{k}}\ast I_{\alpha}^{F_{k}}$ through element-wise multiplication.
@@
$I_{k}=\left(1-I_{\alpha}^{F_{k}}\right)\cdot I_{k-1}+I_{\alpha}^{F_{k}}\ast I_{c}^{F_{k}},\ k\in[K]$ (1)
@@
where $F_{k}$ denotes the $k$-th foreground. The final composite image $I_{K}$ is gained after $K$ iterations. This mathematical formulation explicitly captures two fundamental mechanisms: (1) The multiplicative attenuation of background visibility through successive transparency operations; (2) The nested modulation of foreground layer contributions based on depth-ordered occlusions.
3.2 PSDiffusion